# Observing network dynamics through sentinel nodes

### v1

@@ II.2 Characteristics of good node sets
Our first observation is that nodes in $S$ tend to have a degree sequence that is balanced across the average degree in the entire network. We show this pattern in Fig. 3a for the dolphin network. The top panel of Fig. 3a shows the degree histogram of all nodes in the network. The degree distribution is bimodal, with a predominance of nodes with degree 1 and a secondary group of nodes with degree near the mean degree $\overline{k}=5.13$. The lower panels of Fig. 3a show the aggregate degree histograms realized by the nodes in $S$ obtained from 100 independent runs of the algorithm with $n=1$, $2$, $3$, and $4$, with one panel corresponding to an $n$ value. The degrees of the chosen node are sorted such that the blue bars represent the smallest degree node in the optimized node set in the 100 runs, orange the second smallest, red the third smallest, and light blue the largest degree nodes. We find that the algorithm balances the degrees of the selected nodes such that both small-degree and large-degree nodes tend to exist in a node set when $n\in\{3,4\}$ (and to a lesser extent when $n=2$). For example, in Fig. 1f, the $n=4$ nodes selected for $S$ have degrees 1, 6, 7, and 8. The $x_{i}^{*}$ values for these four nodes, shown by the four blue lines in Fig. 1f, appear above and below the target values, $\overline{x}$, depending on whether the degree of the node is large and small, respectively. As a result, the average of these four lines provides an accurate approximation to $\overline{x}$. Qualitatively the same phenomenon also occurs for $n=2$ and $n=3$, as shown in Fig. 1d and e, respectively.
Figure 3: Degree histogram of the original network and that from the nodes in the optimized node set. (a) Dolphin network. (b) BA network. The original degree histograms for the two networks are shown in the top panels of (a) and (b). We show the degree histograms of 100 optimized node sets with different sizes, $n$, in the second to the bottom panels in each of (a) and (b). In these panels, the color of the bar represents the order of the node in $S$ in terms of the degree (blue: smallest-degree node, orange: second smallest-degree node, red: third node, light blue: fourth node, green: fifth node, yellow: sixth node). The inset in (b) shows the count of nodes with $k_{i}>20$.
We show in Fig. 3b the results of the same analysis for the BA network used in Fig. 2b. The network has hub nodes, i.e., nodes with large degrees. The top hub nodes, with $k_{i}=$ 38, 40, 66, 76, and 81, are indicated by discrete peaks in the inset of the top panel of Fig. 3b. The bottom panels of Fig. 3b indicate that our optimization algorithm does not select these hub nodes in any of the 100 runs for each $n\in\{1,2,3,6\}$. Instead, the algorithm selects nodes the degree of which is close to the mean; note that the mean degree is close to the minimum degree in BA networks. More specifically, the algorithm selects a mixture of nodes with small and moderately large degrees, but avoiding nodes with larger degrees.
@@
The degree histograms of the optimized node sets shown in Fig. 3 appear to markedly differ from that of the entire network. To quantify the difference, we computed the Kullback-Leibler divergence, $D_{\rm KL}$, between the optimized node sets and the original network in terms of the degree distribution (see SI section S5 for the methods).
We show the $D_{\rm KL}$ values for the dolphin network by the blue circles in Fig. 4a. For comparison, the green squares represent the average $D_{\rm KL}$ between the original degree distribution and completely random degree distributions constructed with the same number of nodes (i.e., $100n$ nodes) as that used for constructing the degree distribution from the $100$ optimized node sets. Figure 4a indicates that the discrepancy between the optimized node sets and the original networks in terms of the degree distribution is substantially large when $n$ is small; note the logarithmic scale on the vertical axis. As $n$ becomes larger, both the optimized and the random node sets tend to become more similar to the original network in terms of the degree distribution. However, the degree distribution of the optimized node sets is significantly different from the original degree distribution at all $n\in\{1,\ldots,12\}$. Therefore, although our method tends to select a mix of large- and small- degree nodes, it does not sample the degree values uniformly at random even for relatively large node sets (i.e., up to $n=12$). We have verified that the optimized node sets have smaller approximation errors than completely random node sets of equivalent size for the same range of $n$ (see SI section S2), extending the results for the dolphin network shown in Fig. 2a for $n=4$. It should be noted that the approximation error for both optimized and completely random node sets decreases as $n$ increases (see SI section S2), which is expected.
@@
Figure 4: Comparison between the degree distribution for the optimized node sets and the completely random node sets on two networks. We show the Kullback-Leibler divergence, $D_{\rm KL}$, to the degree distribution of the original network. The blue circles represent the $D_{\rm KL}$ values for the optimized node set. The green squares represent the average $D_{\rm KL}$ over 100 sets of simulations with 100 completely random node sets each; the 95% confidence interval for each mean is smaller than the square marker and therefore not shown. (a) Dolphin network. (b) BA network.
We have also found that the average degree of the nearest neighbors of the nodes in the optimized sentinel node set, sentinel nodes’ local clustering coefficients, and their community membership systematically deviate from these quantities for uniformly randomly selected nodes. However, heuristically constructed node sets exploiting the observed network properties, such as the avoidance of hub nodes (informed by Fig. 3) or multiple nodes from the same community, only marginally, though significantly, decrease the approximation error, $\varepsilon$, compared to completely random node sets. See SI section S6 for these results. Therefore, we conclude that running our optimization algorithm is a necessary step to realize a high accuracy in approximating $\overline{x}$.
@@
II.3 Transfer learning: Effectiveness of optimized node sets across different dynamics
@@
Our optimization algorithm does not use just the information on the network structure but also implicitly requires the dynamic model running on the networks. The challenge is that, in practice, we often do not know the actual dynamics that produce the observed data. Therefore, here we ask whether and to what extent our optimization algorithm enables transfer learning, that is, the application of node sets optimized with one dynamics, called the training dynamics, to another dynamics, called the test dynamics.
@@
We again use the dolphin network as an example. In each panel of Fig. 5a, we plot the approximation error, $\varepsilon$, of each optimized (blue circles), degree-preserving (orange triangles), and completely random (green squares) node set. For each node set, we show the approximation error for the training dynamics on the horizontal axis and the test dynamics on the vertical axis. We also show the average result (large symbols), as obtained for 100 node sets of each type.
As expected, node sets optimized on the training dynamics have consistently smaller $\varepsilon$ when evaluated on the same dynamics than on an alternative test dynamics. However, although $\varepsilon$ has more variance when evaluated on the test dynamics, the node sets optimized under a different training dynamics still outperform the completely random node sets on average (see Fig. 5a). See Fig. 5b–d for an example. In nine out of the twelve pairs of training and test dynamics, the optimized node sets also outperform the degree-preserving node sets (see Fig. 5a). Therefore, the optimized node sets are transferrable between different dynamics to some extent. This result is striking in particular because different dynamics may have different bifurcation structures. For example, the coupled double-well dynamics have saddle-node bifurcations, and the SIS dynamics have transcritical bifurcations. Additionally, the completely random node sets in each panel show an approximately linear relationship. Therefore, the random node sets that happen by chance to be good sentinels for a specific training dynamics continue to be good sentinels for the test dynamics on average. This result further supports our claim that the performance of node sets is transferrable across different dynamics.
@@
Figure 5: Transfer learnability of the sentinel node approximation. (a) Approximation error for optimized node sets evaluated on training and test dynamics. Using the dolphin network, we computed 100 optimized node sets on one (training) dynamics. For the optimized node sets, we show the approximation error when evaluated on the same training dynamics, shown by the blue circles on the horizontal axis, and on an alternative (i.e., test) dynamics, shown on the vertical axis. We also show the approximation errors of 100 degree-preserving node sets (orange triangles) and 100 completely random node sets (green squares) evaluated on the training and test dynamics. The large markers show the average approximation error over the 100 node sets for each type of node set. (b) Dolphin network. The large blue nodes indicate the members of the optimized node set with $n=4$ nodes that attains the lowest approximation error on the coupled double-well training dynamics. (c) Sentinel node approximation for the coupled double-well dynamics on the dolphin network. The states of the nodes marked in (b) are shown in blue. The states of the remaining nodes are shown in gray. The black line represents the unweighted average state of all nodes, $\overline{x}$. The purple line represents the sentinel node approximation of the node set marked in blue. (d) Sentinel node approximation for the SIS dynamics on the same network when the node set is optimized for the coupled double-well dynamics. The blue lines show $x_{i}^{*}$ for the nodes marked in (b).
@@
As before, we assess the generalizability of these observations with an ANOVA. The dependent variable is $\ln\varepsilon$ evaluated on the given test dynamics. The independent variables are the training dynamics, test dynamics, network, and node set type (see SI section S7 for details). Despite the substantial increase in $\varepsilon$ due to the lack of knowledge about the test dynamics, optimized node sets still achieve $\varepsilon$ that is 5.74 times smaller than completely random node sets ($p<10^{-7}$) and 1.55 times smaller than degree-preserving node sets ($p<10^{-7}$). We obtained similar results on the feasibility of transfer learning in the case of weighted networks (see SI section S9) and directed networks (see SI section S10). In sum, these results suggest that our optimization algorithm can select good sentinel node sets when we do not know the test dynamics, or even the type of bifurcation.
II.4 Weighted averaging
We have restricted ourselves to unweighted averages of $x_{i}^{*}$ as our approximators. Previous work has considered the case $n=N$ and chosen the weight vector $(a_{1},\ldots,a_{N})$ to be either the degree vector (i.e., $a_{i}$ is proportional to the degree of the $i$th node) [9, 23, 15] or a particular eigenvector [10, 11, 13, 24, 14]. In this section, we keep $n$ small and additionally consider optimization of the weights of each node to allow a weighted average of $x_{i}^{*}$ given by $\overline{x}^{\prime}=\sum_{i=1;i\in S}^{N}a_{i}x_{i,\ell}^{*}$, subject to $\sum_{i=1;i\in S}^{N}a_{i}=1$ and $a_{i}\geq 0$ $\forall i\in S$. We find these weights using a quadratic programming step (see the Methods section for details) inserted into our original algorithm.

### v3

@@ II.3 Characteristics of good sentinel node sets
Our first observation is that nodes in $S$ tend to have a degree sequence that is balanced across the average degree in the entire network. We show this pattern, as observed in the dolphin network, in Fig. 3. In Fig. 3a, we show the degree histogram of all nodes in the network. We observe a bimodal degree distribution, with a dominant peak of nodes with degree $1$ and a secondary group of nodes scattered around the mean degree $\overline{k}=5.13$. In Fig. 3b-e, we show the aggregate degree histograms realized by the nodes in $S$ obtained from $100$ independent runs of the algorithm with $n=1$, 22, 33, and 44—one panel corresponding to each $n$ value. The degrees of the chosen nodes are sorted such that the blue bars represent the smallest degree node in the optimized node set over the $100$ runs, orange the second smallest, red the third smallest, and light blue the largest degree nodes. We find that the algorithm balances the degrees of the selected nodes such that both small-degree and large-degree nodes tend to exist in the sets. This is quite clearly observed when $n\in\{3,4\}$, and to a lesser extent when $n=2$.
The reason is that in order to capture the behavior of the average $\overline{x}$, we must balance nodes from different groups in the network. For example, in Fig. 1f, the $n=4$ nodes selected for $S$ have degrees 1, 6, 7, and 8. The $x_{i}^{*}$ values for these four nodes, shown by the four blue lines, illustrate precisely this balance: some appear above $\overline{x}$ and some below, depending on whether their degrees are large or small. As a result, the average over these four selected nodes offers an accurate approximation of $\overline{x}$. Qualitatively, we observe the same effect also for $n=2$ and $3$, as shown in Fig. 1d and e, respectively.
In Fig. 3f-j, we repeat the same analysis for a Barabási-Albert (BA) network with $N=10^{3}$ nodes. Here, the largest hub nodes, with $k_{i}=38,40,66,76$, and $81$, appear as discrete peaks in the inset of Fig. 3f. By examining the optimized sentinel node sets in Fig. 3g-j, we find that this inset is empty, implying that our optimization algorithm does not select these hub nodes under all values of $n\in\{1,2,3,6\}$, in any of the $100$ runs. Instead, we observe that $S$ comprises a mixture of nodes with both small and moderately large degrees in all cases, however, avoiding the largest hubs. This result further strengthens our observation above, pertaining to individual nodes, that $S$ tends to exclude the hubs (Supplementary Section S7; see Fig. S14).
To complete our analysis, in Supplementary Section S7 (see Fig. S8), we repeat this process for each of our $20$ networks. We find that the effect is quite consistent: sentinel node sets tend to probe the network’s degree sequence and select nodes with diverse degrees, both above and below the average. At the same time, however, they typically lack the largest hubs, remaining within the bulk of the degree distribution and avoiding its extreme values.
@@
As a result, the degree histograms of the optimized node sets appear to markedly differ from that of the entire network (see Fig. 3). To quantify this difference, we compute the Kullback-Leibler divergence, $D_{\rm KL}$, between the optimized node sets and the original network’s degree distribution (see Supplementary Section S9). In Fig. 4a, we show $D_{\rm KL}$ for various values of $n$ between the original network degree distribution and that of the $100$ optimizes sets (blue circles). We also show the results obtained from completely random node selections (i.e., $100n$ nodes; green squares).
For small $n$, the sentinel degree distribution features a discrepancy from that of the entire network, as observed through its much larger $D_{\rm KL}$ than the completely random samples; note the logarithmic scale on the vertical axis. Therefore, $S$ indeed does not simply sample nodes that are representative of the network’s degree sequence. As $n$ is increased, both the optimized and the completely random node sets approach the network’s degree distribution. However, throughout the entire range of examined $n$ values ($n\in\{1,\ldots,12\}$), the degree distribution of the optimized node sets remains significantly different from that of the original network while maintaining a smaller $\varepsilon$ (see Supplementary Section S10). Therefore, although our method tends to select a mixture of large and small nodes, it does not sample the degree values uniformly at random even for relatively large node sets. Similar results are also obtained for our BA network (Fig. 4b).
@@
Figure 4: Degree distribution of optimized vs. random sets. We collected $100$ optimized sets of $n=1,\dots,12$ nodes from the dolphin network and extracted their collective degree distributions. (a) The Kullback-Leibler divergence, $D_{\rm KL}$, between the optimized sets and the original network as $n$ varies (blue circles). The results obtained from $100$ completely random node sets are also shown (green squares); the 95% confidence interval for each mean is smaller than the symbols and therefore not shown. For all $n$ values, and especially for the smaller sets, we find that the sentinel $D_{\rm KL}$ is significantly larger than that of the random sets. (b) Similar results are obtained for the BA network. Here the gap between the two set types is even more pronounced, likely due to the presence of hubs in the network and their absence in $S$.
In Supplementary Section S11, we consider several additional node set features. The first six capture the previously considered features, i.e., nearest neighbor degree, clustering etc., only this time averaged over the $n$ nodes in the set. Beyond these set averages, we also consider two additional inherent node set features: (1) the maximum number of nodes in a node set which come from the same network community, and (2) the average pairwise distance between nodes in a node set (Supplementary Section S11). We found, using random forest models, that these node set features combined can, indeed, explain $\varepsilon$ to a large extent (see Table S17). In particular, the set’s average degree should be neither too small nor too large for suppressing $\varepsilon$ (see Fig. S18). We emphasize that this result does not imply that all the $n$ nodes should have intermediate degrees. As we have shown above, it is generally better for the $n$ nodes to cover a range of degree while avoiding the extremes.
With these observations, we can now construct heuristic algorithms that bypass our brute-force optimization method. Such algorithms determine sentinel node sets based on the observed sentinelity features above. For example, one can prioritize node sets that have the desired degree sequences, as motivated by Fig. 3, or select nodes to ensure that they are sampled from distinct network communities. We examine these heuristic algorithms in Supplementary Section S11. We find that, while these algorithms significantly reduce $\varepsilon$, they are still substantially inferior to our optimized node sets (see Table S18). Therefore, to achieve sufficient accuracy in retrieving the global state $\overline{x}$, one must still run our full optimization algorithm.
@@
II.4 Transfer learning: Effectiveness of optimized node sets across different dynamics
@@
Our optimization algorithm relies not only on network structure but also implicitly on the dynamics running on the network. The challenge, however, is that in practice, the actual dynamics governing the observed data are often unknown. This raises the question: to what extent does our optimization algorithm support transfer learning? In other words, can node sets optimized under one dynamics—referred to as the training dynamics—be effectively applied to a different dynamics, i.e. the test dynamics?
@@
We begin, once again, by examining the dolphin network, considering all potential pairings of training and test dynamics. Note that here we also use mismatched pairs of the dynamics and network (e.g., the gene-regulatory dynamics on the dolphin network) to test the possibility of transfer learning. In Fig. 5a, each panel displays the approximation error, $\varepsilon$, as obtained from all our optimized (blue circles), degree-preserving (orange triangles), and completely random (green squares) node sets. For each node set, we show the approximation error for the training dynamics on the horizontal axis and the test dynamics on the vertical axis. We also present the average result (large symbols), computed from the $100$ node sets of each type.
As expected, node sets optimized on a specific training dynamics have consistently smaller $\varepsilon$ when evaluated on the same dynamics than on an alternative test dynamics. Indeed, they were optimized for that specific dynamics. For example, in Fig. 5a7, we used the double-well dynamics to train the model and obtain the set $S$, and then tested $S$’s performance against the susceptible-infectious-susceptible (SIS) dynamics. When we use the SIS dynamics as the training dynamics, the blue circles condense around $\varepsilon=10^{-5}$ (the horizontal axis of Fig. 5a2, a5, and a12) indicating an almost perfect prediction. However, on the vertical axis of Fig 5a7, capturing the extent of transfer learning from the double-well to SIS, we observe an average error of $\varepsilon=5\times 10^{-3}$. We also witness a broader spread across our $100$ trials, together indicating that the transfer from the training dynamics to the test dynamics is imperfect.
The crucial point is, however, that across all training-test pairings, the optimized node sets, on average, outperform the random node sets (blue circles vs. green squares). This result indicates that a set optimized under one dynamics can indeed be useful for predicting the state of another dynamics. We also tested degree-preserving node sets, finding that our optimized sentinels continue to exceed their performance in ten out of the twelve tested pairings. Therefore, the optimized node sets are, to some extent, transferable between different dynamics. This result is striking in particular because different dynamics may have different bifurcation structures. For example, the coupled double-well dynamics have saddle-node bifurcations, whereas the SIS dynamics have transcritical bifurcations. Still, as Fig. 5a7 indicates, node sets trained on one dynamics remain predictive on the other. To demonstrate this explicitly, in Fig. 5b, we show a specific sentinel nodes set, optimized on the dolphin network under the double-well dynamics (blue nodes). Indeed, the set is shown to successfully capture the transition pattern of its training dynamics (Fig. 5c). This positive result is expected. However, in Fig. 5d, we use the same sentinel node set to track the SIS dynamics, for which it was never trained. The figure shows that the learning against the double-well dynamics successfully transfers to the SIS dynamics.
The root of this transferability is likely due to the fact that sentinelity is primarily a structural characteristic. Therefore, it only weakly depends on the specific dynamics, and mainly driven by the network topology. This insight is crucial when we have little knowledge of the network dynamics. In this case, we can train $S$ against a presumed dynamics and then use the obtained sentinel node set to observe the activity of the hidden dynamical system. Interestingly, the completely random node sets in each panel (green circles in Fig. 5a) show an approximately linear relationship. This result indicates that random node sets that happen by chance to be good sentinels for a specific training dynamics continue, on average, to be good sentinels also for the test dynamics. This observation further supports the notion that sentinelity is embedded in the topology and hence it is transferable across different dynamics.
In Fig. 5, we analyzed transferability on the relatively small dolphin network. This choice allowed us to track the specific node sets and visualize them, as we did in Fig. 5b. To complete this analysis, in Supplementary Section S12, we repeat the same experiment on a large financial network, comprising $N\approx 8\times 10^{4}$ nodes, thus covering a range of scales (Fig. S19). In Supplementary Section S12, we also show similar transferability results for a mobility-based human network (Fig. S20) and an airport network (Fig. S21). The positive results obtained suggest potential relevance of the proposed method to real epidemic outbreaks.
@@
Figure 5: Transfer learnability of the sentinel node approximation. We used the dolphin network and our four dynamic models to construct all potential pairs of training and test dynamics. (a) The approximation error $\varepsilon$ obtained from the test dynamics vs. that of the training dynamics for our optimized node sets (blue), degree-preserving node sets (orange), and completely random node sets (green). For example, in panel (a1), we extract the sentinels using the mutualistic species dynamics ($x$-axis, training) then examine their performance against the coupled double-well dynamics ($y$-axis, test). For each node selection method, we present results from $100$ independent sets by small symbols and the averaged $\varepsilon$ by large symbols. In all but two panels, we find that our optimized node sets perform best, both against the training dynamics and, most crucially, against the test dynamics. (b) Dolphin network. The large blue nodes indicate the members of the optimized node set with $n=4$ nodes that attains the lowest approximation error on the coupled double-well training dynamics. (c) Sentinel node approximation for the coupled double-well dynamics on the dolphin network. The states of the nodes marked in (b) are shown in blue. The states of the remaining nodes are shown in gray. The average state $\overline{x}$ (black solid line) and the sentinel node approximation (purple solid line) are also shown. As expected, the sentinel node approximation follows $\overline{x}$ quite tightly. (d) Sentinel node approximation for the SIS dynamics on the same network, using the node set optimized for the coupled double-well dynamics. Despite being trained on a different dynamics, the four blue nodes still offer a reliable approximation for the SIS dynamics.
@@
As before, we assess the generalizability of these observations with an ANOVA. The dependent variable is $\ln\varepsilon$ evaluated on the given test dynamics. The independent variables are the training dynamics, test dynamics, network, and node set type (see Supplementary Section S12 for details). Despite the substantial increase in $\varepsilon$ due to the lack of knowledge about the test dynamics, optimized node sets still achieve $\varepsilon$ that is 7.43 times smaller than completely random node sets ($p<10^{-7}$) and 1.52 times smaller than degree-preserving node sets ($p<10^{-7}$) (see Table S19). We obtained similar results on the feasibility of transfer learning in the case of weighted networks (Table S3) and directed networks (Table S5). In summary, these results suggest that our optimization algorithm can select good sentinel node sets when we do not know the test dynamics, or even the type of bifurcation.
We emphasize that these results in no way imply that the the observed network behavior is independent of the dynamics. Clearly, the network activity $\overline{x}$, which we wish to predict, and its bifurcation patterns, all arise from the interplay between structure and dynamics. Therefore, we expect that under different dynamics the system’s behavior may be fundamentally distinct. This phenomenon is clearly observed in Fig. 5c,d, where the same network hosts qualitatively different behavior when coupled with different dynamics (double-well vs. SIS). The point is that the sentinels that can be used to observe these distinct $\overline{x}$ curves do transfer across dynamics, because they are mainly determined by topology and at large independent of the system’s dynamics.
In the case of undirected and unweighted networks, we found that degree-preserving random node sets sometimes perform better than optimized node sets in terms of transfer learning. This phenomenon was observed specifically on large networks, or when our gene-regulatory dynamics was used either for the training or for the test dynamics (see Fig. S22). This observation can offer guidelines for further strengthening the transfer learning. For example, given a large network, one may benefit by running our optimization algorithm on one dynamics to determine sentinel node sets. Then, by extracting each set’s degree sequence, one can draw a degree-preserving node set, thus combining the strength of both our optimization method and the advantages of degree preservation.
In real systems, we often encounter uncertainties such as hidden nodes or missing edges. Therefore, we systematically examine the robustness of our transfer learning against such missing network components. We find that even if one constructs sentinel node sets with $\approx 10\%$ of the nodes/links unobserved, the obtained optimized sets are still more accurate then random ones at approximating $\overline{x}$ for the full network, i.e., where all components are present (Supplementary Section S13). Optimized node sets are also transferable under uncertainties in the coupling strength $D$ when only some part of the network undergoes changes in $D$ (Supplementary Section S14).
II.5 Sentinel nodes in empirical time series data
In empirical complex systems, the dynamics generating the data at each node is usually unknown. The underlying network upon which the dynamics takes place is also subject to uncertainties and is often only partially mapped. Furthermore, the dynamics may not be in equilibrium, and hence the measured $\overline{x}$ may not properly represent the system’s equilibrium. To assess the applicability of our sentinel node approximation under these challenging conditions, we collected multivariate time series data from the brain activity of healthy human individuals using functional magnetic resonance imaging (fMRI). The data, extracted from the Human Connectome Project [25], cover $1,003$ participants, each tracked at a spatial resolution of $N=300$ nodes. An explicit and established map of the network structure is lacking in these data. Therefore, for each participant, we used the first $4,700$ time points to estimate the network structure, and the remaining $100$ time points for recording the node activities $x_{i}$ (Supplementary Section S15).
The second challenge is that, as noted above, we do not have the specific dynamics for this system. Fortunately, this is precisely the scenario where transfer learning is designed to help, allowing us to obtain $S$ under one dynamics and approximate $\overline{x}$ under another. Therefore, we implemented the coupled double-well dynamics on the reconstructed brain network, and simulated the state of the system over the same range of $D$ values as previously used (Fig. 1). We then extracted the optimized sentinel node sets and tested their prediction against the observed brain time series data. To assess the performance of our optimized node sets, we measured the error $\varepsilon$ and compared it to that obtained from completely random node sets (Fig. 6; see Supplementary Section S15, including Tables S20 and S21, for statistical results). We find that, under the optimized sentinel node sets, $\varepsilon$ is on average $6.09\%$ smaller than under the completely random node sets. With degree-preserving node sets, $\varepsilon$ is on average $3.41\%$ smaller than that of random sets. These results were quantitatively similar when we determined optimized node sets using any of the other three dynamics or a network variant of the Wilson-Cowan dynamics for neural masses [10] (see Supplementary Section S15). At first glance, these improvements may seem modest, but to truly appreciate their significance, we must consider the conditions under which they were achieved. First, fMRI data are relatively noisy. Second, our attempt is to predict $\overline{x}$ without any a priori data on the network or the relevant dynamics. And yet, our method was still able to detect nodes with a discernible predictive advantage. In fact, we have shown that the results do not notably depend on the dynamics. Furthermore, there is no reason to assume that the brain dynamics adhere to any of our four considered model dynamics or the Wilson-Cowan dynamics. Therefore, even this seemingly small enhancement in accuracy is sufficient to demonstrate the potential practical utility of our transfer learning in empirical environments.
Figure 6: Sentinel node sets in empirical data. We tested our optimization algorithm on multivariate time series fMRI data recorded from $1,003$ healthy human individuals. We estimated networks from a training set of each individual’s time series data, simulated the coupled double-well dynamics, and then selected 100 sentinel node sets based on the simulation results. The distribution of the average approximation error $\varepsilon$ of each individual’s 100 sentinel node sets, across all the individuals, is shown in blue; the corresponding distributions for degree-preserving and completely random node sets are shown in orange and green, respectively. The mean of each distribution is indicated by the vertical purple lines.
II.6 Weighted averaging
Up to this point, we restricted ourselves to unweighted averaging over the sentinel node set, in which all node activities within $S$ add uniformly towards the approximation of $\overline{x}$. More generally, we may relax this restriction and assign non-uniform weights to all sentinels, approximating the system’s state via $\overline{x}^{\prime}=\sum_{i=1;i\in S}^{N}a_{i}x_{i,\ell}^{*}$. Here, the weights $a_{i}$ are subject to $\sum_{i=1;i\in S}^{N}a_{i}=1$ and $a_{i}\geq 0$ $\forall i\in S$, i.e. positive and normalized. This setting provides an additional degree of freedom to our optimization, which can potentially improve the approximation accuracy. To optimize the weights, we add a quadratic programming step to our original algorithm (see the Methods Section).
Such weighted averages are inspired by existing methods, such as GBB or DART, that have indicated that, to predict the system state, it is optimal to assign different weights to all nodes. For example, nodes are weighted by their degree ($a_{i}=k_{i}/\sum_{\ell=1}^{N}k_{\ell}$) in GBB [9, 26, 15] and by the $i$th entry of the principal eigenvector in DART [10, 11, 13, 27, 14]. The main difference is that in these applications one considers all nodes ($n=N$) and assignes the weights based on theoritically motivated heuristics. In contrast, in our application, we continue to demand $n\ll N$ and tune the weights using machine learning tools.